# CAKE: Cascading and Adaptive KV Cache Eviction with Layer Preferences

### v1

@@ Appendix B Limitation and Future Work
CAKE introduces a novel perspective on KV cache management by considering the model’s varying attention patterns across layers and comprehensively measuring layer-specific preferences for KV cache allocation. This approach enables effective adaptation to different models and input contexts, optimizing the use of limited memory resources. Despite these strengths, several limitations and areas for future work remain. While CAKE improves cache management at the layer level, it does not address finer-grained dynamics within layers. Future work could explore integrating head-level attention patterns with layer-specific variations, potentially offering more nuanced control over cache allocation. Additionally, although we demonstrate CAKE’s compatibility with existing eviction methods (Section 5.5) and quantization methods (Appendix 5.5), there is potential to further enhance memory efficiency by combining it with other KV cache optimization techniques such as cache merging (Liu et al., 2024b) and pruning (Xu et al., 2024). This integration could significantly expand CAKE’s applicability and resource-saving capabilities. In the future, we will continue to explore strategies for efficient LLMs, focusing on both memory efficiency and computational optimization to further push the boundaries of resource-constrained inference.
@@ Appendix C More Implementation Details
In this section, we provide more detailed implementation information for CAKE. Our algorithm is divided into two phases: prompt prefilling and token decoding. During prompt prefilling, we utilize preference-guided cascading cache management (Section 4.2) based on the preference-prioritized adaptive allocation strategy (Section 4.1) to allocate appropriate cache sizes for different layers. During preference-guided cascading cache management, we manage the KV cache by updating it using eviction operation based on the attention-shift tolerant eviction indicator (Section 4.3). For the preference-prioritized adaptive allocation strategy, the temperature parameters $\tau_{1}$ and $\tau_{2}$ are set to $0.2\sim 2$ and $0.4\sim 3$ respectively, optimized through grid search. For our experimental models, we aggregate each quantized value to represent its layer characteristics. For the attention-shift tolerant eviction indicator, we fix $\gamma=200$. Following SnapKV (Li et al., 2024b), we use an observation window of size $S_{w}=32$ and utilize a pooling layer clustering information. In the decoding stage, once the attention mechanisms for layers have been well established, we fix the budget size for each layer according to the allocation results obtained from the preference-prioritized adaptive allocation. Listing 1 provides PyTorch-style pseudo-code for CAKE’s implementation with FlashAttention.
@@ Appendix E Comparison and Compatibility with Quantization Methods
In this section, we evaluate CAKE against two quantization approaches: KIVI (Liu et al., 2024e), a state-of-the-art method using asymmetric KCVT quantization (quantizes Key cache per-channel and Value cache per-token), and a computationally efficient channel-wise quantization variant for both key and value caches (denoted as KCVC). As is shown in Figure 9, under matched compression ratios, CAKE outperforms full cache (33.23 vs. 33.07) and both quantization methods at 25% and 12.5% compression ratios. Notably, when combining KIVI-INT4 with CAKE, it demonstrates significant advantages in maintaining performance at high compression ratios. For instance, KIVI-INT4 with CAKE achieves 32.51 at 12.5% compression ratio, outperforming KIVI-INT2 (32.17) while using the same storage overhead. Even at a very aggressive 6.25% compression ratio, the combined approach still maintains strong performance (32.48), showing the effectiveness of integrating eviction and quantization strategies.
@@ Appendix F Extended Experimental Results on LongBench
In this section, we provide comprehensive experimental results on LongBench (Bai et al., 2023), a benchmark focused on long-context understanding with input lengths ranging from 1,235 to 18,409 tokens. We present detailed performance evaluations across cache budgets from 64L to 2048L for three base models: Llama2-7B-Chat (Touvron et al., 2023), Llama3.1-8B-Instruct (Dubey et al., 2024), and Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) (Appendix F.1). To demonstrate CAKE’s generalizability, we conduct additional experiments on two more LLM architectures: Qwen2.5-7B-Instruct (Team, 2024) and Gemma-7B-Instruct (Team et al., 2024) (Appendix F.2). Furthermore, we evaluate CAKE’s scalability on larger models ranging from 13B to 70B parameters, including Llama2-13B-Chat, Qwen2.5-32B-Instruct, and Llama3-70B-Instruct (Appendix F.3).
@@ F.1 Detailed Performance Analysis Across Cache Budgets
Table 6: Performance comparison over 16 datasets of LongBench on Llama3.1-8B-Instruct for cache budgets from $64L$ to $2048L$ The best result is highlighted in bold, the second best in underline.
@@ Appendix G Extended Experimental Results on Needlebench
In this section, we provide more detailed experimental results on NeedleBench (Li et al., 2024a), using Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) and Llama3.1-8B-Instruct (Dubey et al., 2024) as backbone LLMs. We examine three distinct tasks on NeedleBench: (1) Single-Needle Retrieval (S-RT): Assessing precision in locating individual details within extensive texts. (2) Multi-Needle Retrieval (M-RT): Evaluating the retrieval of multiple related information pieces dispersed throughout lengthy texts. (3) Multi-Needle Reasoning (M-RS): Measuring complex reasoning abilities by extracting multiple information elements from extended contexts to answer questions.
@@
Prefill+Decoding
Length Method Overall Generation Time (s) Prompt Prefill Time (s) Decoding Time (s) Decoding Time per Token (ms)
@@ Appendix H Additional Experiments and Analysis on Efficiency
In this section, we present a detailed time breakdown during both prompt prefilling and token decoding to better assess CAKE’s effectiveness across different inference stages compared to other KV eviction methods, including SnapKV and PyramidKV with various allocation strategies. All methods are implemented as extensions of FlashAttention-2 (Dao, 2023) using Mistral-7B-Instruct-v0.3, and their performance was benchmarked against a full cache baseline utilizing FlashAttention. To measure computational time more precisely throughout the entire text generation phase, we have separately evaluated the prompt prefilling time and the duration of autoregressive decoding. By fixing the generation token count at 1024, the experiment simulates typical long-context processing scenarios in LLMs, which are characterized by lengthy inputs and concise outputs.
As shown in Table G, while dramatically lowering latency in the decoding phase, CAKE, SnapKV, and PyramidKV achieve comparable prompt prefilling times. For SnapKV and PyramidKV with fixed allocation strategies, the prompt prefilling stage requires waiting for KV caching and performing one eviction operation per layer, resulting in $L$ eviction operations. However, compared to the computation-intensive prefill stage, the eviction operation is relatively inexpensive and can be considered negligible. Notably, while CAKE utilizes preference-guided cascading cache management, it achieves similar timing to SnapKV and PyramidKV with fixed allocation strategies. This is attributed to the parallelizable execution of KV cache eviction between layers (Algorithm 1, lines 9–14), which allows the dynamic update process to be incorporated into a single eviction operation. Therefore, the total execution time ultimately equates to that of performing $L$ eviction operations.
@@ I.2 Influence of Temperature Parameter
(e) Llama3.1-8B-Instruct
(f) Mistral-8B-Instruct

### v2

@@ Appendix B Limitation and Future Work
CAKE introduces a novel perspective on KV cache management by considering the model’s varying attention patterns across layers and comprehensively measuring layer-specific preferences for KV cache allocation. This approach enables effective adaptation to different models and input contexts, optimizing the use of limited memory resources. Despite these strengths, several limitations and areas for future work remain. While CAKE improves cache management at the layer level, it does not address finer-grained dynamics within layers. Future work could explore integrating head-level attention patterns with layer-specific variations, potentially offering more nuanced control over cache allocation. Additionally, although we demonstrate CAKE’s compatibility with existing eviction methods (Section 5.5) and quantization methods (Appendix 5.5), there is potential to further enhance memory efficiency by combining it with other KV cache optimization techniques such as cache merging (liu2024minicache) and pruning (xu2024think). This integration could significantly expand CAKE’s applicability and resource-saving capabilities. In the future, we will continue to explore strategies for efficient LLMs, focusing on both memory efficiency and computational optimization to further push the boundaries of resource-constrained inference.
@@ Appendix C More Implementation Details
In this section, we provide more detailed implementation information for CAKE. Our algorithm is divided into two phases: prompt prefilling and token decoding. During prompt prefilling, we utilize preference-guided cascading cache management (Section 4.2) based on the preference-prioritized adaptive allocation strategy (Section 4.1) to allocate appropriate cache sizes for different layers. During preference-guided cascading cache management, we manage the KV cache by updating it using eviction operation based on the attention-shift tolerant eviction indicator (Section 4.3). For the preference-prioritized adaptive allocation strategy, the temperature parameters $\tau_{1}$ and $\tau_{2}$ are set to $0.2\sim 2$ and $0.4\sim 3$ respectively, optimized through grid search. For our experimental models, we aggregate each quantized value to represent its layer characteristics. For the attention-shift tolerant eviction indicator, we fix $\gamma=200$. Following SnapKV (li2024snapkv), we use an observation window of size $S_{w}=32$ and utilize a pooling layer clustering information. In the decoding stage, once the attention mechanisms for layers have been well established, we fix the budget size for each layer according to the allocation results obtained from the preference-prioritized adaptive allocation. Listing 1 provides PyTorch-style pseudo-code for CAKE’s implementation with FlashAttention.
@@ Appendix E Comparison and Compatibility with Quantization Methods
In this section, we evaluate CAKE against two quantization approaches: KIVI (liu2024kivi), a state-of-the-art method using asymmetric KCVT quantization (quantizes Key cache per-channel and Value cache per-token), and a computationally efficient channel-wise quantization variant for both key and value caches (denoted as KCVC). As is shown in Figure 9, under matched compression ratios, CAKE outperforms full cache (33.23 vs. 33.07) and both quantization methods at 25% and 12.5% compression ratios. Notably, when combining KIVI-INT4 with CAKE, it demonstrates significant advantages in maintaining performance at high compression ratios. For instance, KIVI-INT4 with CAKE achieves 32.51 at 12.5% compression ratio, outperforming KIVI-INT2 (32.17) while using the same storage overhead. Even at a very aggressive 6.25% compression ratio, the combined approach still maintains strong performance (32.48), showing the effectiveness of integrating eviction and quantization strategies.
@@ Appendix F Extended Experimental Results on LongBench
In this section, we provide comprehensive experimental results on LongBench (bai2023longbench), a benchmark focused on long-context understanding with input lengths ranging from 1,235 to 18,409 tokens. We present detailed performance evaluations across cache budgets from 64L to 2048L for three base models: Llama2-7B-Chat (touvron2023llama), Llama3.1-8B-Instruct (dubey2024llama), and Mistral-7B-Instruct-v0.3 (jiang2023mistral) (Appendix F.1). To demonstrate CAKE’s generalizability, we conduct additional experiments on two more LLM architectures: Qwen2.5-7B-Instruct (qwen2.5) and Gemma-7B-Instruct (team2024gemma) (Appendix F.2). Furthermore, we evaluate CAKE’s scalability on larger models ranging from 13B to 70B parameters, including Llama2-13B-Chat, Qwen2.5-32B-Instruct, and Llama3-70B-Instruct (Appendix F.3).
@@ F.1 Detailed Performance Analysis Across Cache Budgets
Table 6: Performance comparison over 16 datasets of LongBench on Llama3.1-8B-Instruct for cache budgets from $64L$ to $2048L$ . The best result is highlighted in bold, the second best in underline.
@@ Appendix G Extended Experimental Results on Needlebench
In this section, we provide more detailed experimental results on NeedleBench (li2024needlebench), using Mistral-7B-Instruct-v0.3 (jiang2023mistral) and Llama3.1-8B-Instruct (dubey2024llama) as backbone LLMs. We examine three distinct tasks on NeedleBench: (1) Single-Needle Retrieval (S-RT): Assessing precision in locating individual details within extensive texts. (2) Multi-Needle Retrieval (M-RT): Evaluating the retrieval of multiple related information pieces dispersed throughout lengthy texts. (3) Multi-Needle Reasoning (M-RS): Measuring complex reasoning abilities by extracting multiple information elements from extended contexts to answer questions.
@@
Prefill+Decoding Length Method Overall Generation Time (s) Prompt Prefill Time (s) Decoding Time (s) Decoding Time per Token (ms)
@@ Appendix H Additional Experiments and Analysis on Efficiency
In this section, we present a detailed time breakdown during both prompt prefilling and token decoding to better assess CAKE’s effectiveness across different inference stages compared to other KV eviction methods, including SnapKV and PyramidKV with various allocation strategies. All methods are implemented as extensions of FlashAttention-2 (dao2023flashattention2fasterattentionbetter) using Mistral-7B-Instruct-v0.3, and their performance was benchmarked against a full cache baseline utilizing FlashAttention. To measure computational time more precisely throughout the entire text generation phase, we have separately evaluated the prompt prefilling time and the duration of autoregressive decoding. By fixing the generation token count at 1024, the experiment simulates typical long-context processing scenarios in LLMs, which are characterized by lengthy inputs and concise outputs.
As shown in Table 13, while dramatically lowering latency in the decoding phase, CAKE, SnapKV, and PyramidKV achieve comparable prompt prefilling times. For SnapKV and PyramidKV with fixed allocation strategies, the prompt prefilling stage requires waiting for KV caching and performing one eviction operation per layer, resulting in $L$ eviction operations. However, compared to the computation-intensive prefill stage, the eviction operation is relatively inexpensive and can be considered negligible. Notably, while CAKE utilizes preference-guided cascading cache management, it achieves similar timing to SnapKV and PyramidKV with fixed allocation strategies. This is attributed to the parallelizable execution of KV cache eviction between layers (Algorithm 1, lines 9–14), which allows the dynamic update process to be incorporated into a single eviction operation. Therefore, the total execution time ultimately equates to that of performing $L$ eviction operations.
@@ I.2 Influence of Temperature Parameter
(a) Llama3.1-8B-Instruct
(b) Mistral-8B-Instruct